# Remote Sensing-Oriented World Model

### v2

@@ 1 Introduction
Figure 1: Given a central input tile, models generate direction-conditioned continuations in four directions. Model A produces richer spatial extrapolations (orange boxes), whereas Model B introduces little new content but still attains a lower FID score (163.4 vs. 206.42), which is misaligned with the world modeling objective. The proposed RSWISE metric (87.3 vs. 56.5) accounts for spatial extrapolation consistency, yielding an evaluation more aligned with the requirements of remote sensing world modeling.
@@
World models have emerged as a frontier in artificial intelligence, showing promise across diverse applications such as robotic navigation (Wu et al., 2023) and autonomous driving (Guan et al., 2025). These models aim to learn the underlying dynamics of environments from limited observations and to predict or reason about unobserved states (Ding et al., 2025). However, most world model studies remain confined to synthetic simulators or constrained scene settings. Synthetic settings lack the complexity and uncertainty of real environments, while constrained scene settings fail to capture reasoning over large spatial structures. As a result, the real-world effectiveness of current world models in spatial reasoning remains largely untested.
@@
Remote sensing provides a uniquely powerful testbed for world models. Satellite and aerial imagery naturally encode ”world-level” structures such as urban road networks (Yu & Fang, 2023), river systems (Tomsett & Leyland, 2019), agricultural mosaics (Khanal et al., 2020), and forest landscapes (Fassnacht et al., 2024). At the same time, high-impact applications—including flood prediction for disaster response (Nguyen et al., 2024) and infrastructure forecasting in urban planning (Wellmann et al., 2020)—require reasoning beyond directly observed regions. Yet, much of remote sensing research has focused on recognition tasks such as classification (Li et al., 2022; Temenos et al., 2023) and semantic segmentation (Sun et al., 2020; Zhang et al., 2023a), leaving the potential of world modeling in this domain unexplored.
@@
This paper bridges these gaps by introducing the first framework for world modeling in remote sensing. We formulate remote sensing world modeling as direction-conditioned spatial extrapolation (defined in the image-grid frame with up, down, left, and right, rather than geographic cardinal directions), where models generate semantically consistent adjacent image tiles given a central observation and directional instruction.
@@
However, introducing world models to the remote sensing domain faces a fundamental evaluation challenge rooted in the limitations of existing assessment paradigms. Current evaluation approaches suffer from critical methodological flaws across two distinct failure modes. Distributional fidelity metrics such as Fréchet Inception Distance (FID) Heusel et al. (2017) measure statistical realism but ignore whether generated tiles follow spatial instructions. As a result, models can obtain deceptively low FID scores by replicating inputs or producing visually plausible but spatially incoherent imagery. Conversely, large language model–based semantic evaluations such as World Knowledge-Informed Semantic Evaluation (WISE) (Niu et al., 2025) capture directional compliance but lack grounding in distributional realism, often rewarding spatially consistent yet environmentally implausible generations.
@@
To overcome these limitations, we introduce RSWISE, the first evaluation framework designed for remote sensing world models. RSWISE integrates distributional fidelity with spatial reasoning consistency through a dual-dimension approach. Specifically, it employs FID to ensure adherence to real-world satellite statistics and leverages GPT-4o to assess whether generated tiles reveal novel yet geographically plausible regions aligned with directional prompts. As shown in Figure 1, RSWISE better reflects geospatial alignment than FID alone, providing a principled basis for fair comparison and progress tracking.
@@
Afterwards, we present RemoteBAGEL, the first remote sensing world model designed for direction-conditioned spatial extrapolation. Instead of producing visually plausible but spatially incoherent completions, RemoteBAGEL explicitly couples generation with spatial reasoning requirements. It is built around two components: (1) a trajectory-based data construction pipeline that transforms raw satellite imagery into instruction-conditioned continuation tasks, and (2) a reconstruction-driven training framework that enforces geographic continuity and semantic coherence during spatial extrapolation. Extensive experiments demonstrate that RemoteBAGEL consistently outperforms state-of-the-art baselines on RSWISE. In summary, our contributions are threefold:
@@ 2.1 World Models and Benchmarks
World models generally divided into two perspectives: models that aim to understand the world by abstracting its underlying mechanisms, and models that aim to predict the future by simulating possible evolutions of the environment (Ding et al., 2025). Early efforts such as World Models (Ha & Schmidhuber, 2018) focused on abstracting the external world to gain a deep understanding of its underlying mechanisms, while subsequent work including PlaNet (Hafner et al., 2019) and the Dreamer family (Hafner et al., 2020; 2022) introduced recurrent state-space models (RSSMs) that designed to facilitate forward prediction purely within the latent space. More recent advances extend this principle into generative modeling: transformer-based models such as TransDreamer (Chen et al., 2024a) and Genie (Bruce et al., 2024), diffusion and VAE-driven approaches for scene extrapolation and controllable driving (Wang et al., 2025; Cai et al., 2023), and the JEPA family (Assran et al., 2023) that reframe world models as self-supervised abstraction learners. Despite this diversity, evaluation remains a central challenge: existing benchmarks such as VBench (Ji et al., 2024), ChronoMagic-Bench (Yuan et al., 2024), TC-Bench (Feng et al., 2024), and WorldModelBench (Li et al., 2025) focus on controlled or synthetic scene settings, but they do not explicitly involve spatial continuity or geospatial semantics at the remote sensing scale in real-world contexts. This gap prevents current evaluations from testing how well world models reason over real-world structures such as rivers, roads, or urban–rural transitions.
@@
VBench Ji et al. (2024) 800 Video FID / Human
WorldModelBench Li et al. (2025) 350 Games / Video Task-specific
ChronoMagic-Bench Yuan et al. (2024) 1649 Video Temporal Consistency
TC-Bench Feng et al. (2024) 150 Video FID
@@ 2.2 Remote Sensing Models
Remote sensing (RS) acquires Earth observations from satellites and aerial platforms, producing imagery that encodes both spectral variation and large-scale spatial structures across diverse environments (Li et al., 2019). Much of RS research has traditionally focused on recognition-oriented tasks, including land-cover classification (Li et al., 2022; Temenos et al., 2023), semantic segmentation (Sun et al., 2020; Zhang et al., 2023a), object detection (Zhang et al., 2023b; Li et al., 2023; Yu & Ji, 2022), and change detection (Bai et al., 2023; Chen et al., 2024b; Li et al., 2024). Building on this recognition paradigm, recent advances in large-scale pretraining have produced RS foundation models such as SkySense (Guo et al., 2024), AnySat (Astruc et al., 2024), SpectralGPT (Hong et al., 2024), and RemoteCLIP (Liu et al., 2024). These models excel at learning transferable representations and support zero-shot inference across downstream tasks (Huo et al., 2025). However, in contrast to world models, current RS methods seldom attempt spatial continuation or reasoning over large geospatial structures.
@@ 3.2 Dataset Curation
To comprehensively evaluate spatial extrapolation under diverse geospatial conditions, we define four representative scenarios that capture complementary challenges in remote sensing world modeling. The first is the general setting of geographic extrapolation, where models extend observed regions into adjacent unseen areas across a broad variety of landscapes. This scenario is designed to incorporate diverse scene types within a single setting, enabling a comprehensive assessment of overall model capability rather than focusing on a specific environment. Beyond this general setting, flood scenarios introduce highly dynamic environmental variations, urban regions emphasize continuity across dense built environments, and rural landscapes highlight consistency within natural and agricultural patterns. These scenarios jointly establish a structured basis for assessing spatial extrapolation across stable, dynamic, structured, and natural contexts.
The benchmark dataset consists of 1,600 curated evaluation instances evenly distributed across these four scenarios: general, flood, urban, and rural. The data are sourced from three publicly available datasets: Sky-SA (Zhu et al., 2025) for general scenes, FloodNet (Rahnemoonfar et al., 2021) for flood events, and LoveDA (Wang et al., 2021) for urban and rural landscapes.
@@
For each scenario, large satellite images are divided into $3\times 3$ overlapping grids of tiles to preserve boundary consistency and spatial autocorrelation. Start tiles are paired with their four cardinal neighbors (up, down, left, right), yielding evaluation triplets. Directional instructions are standardized into fixed prompts (e.g., “Look at what is below this picture”) to ensure fairness across models. Filtering criteria include cloud cover thresholds, resolution consistency, and temporal alignment. A quality assurance process—combining automated checks for artifacts, manual inspection of geographic coherence, and balanced sampling—produces 400 instances per category.
@@ 3.3 RSWISE Evaluation Metrics
where $m$ denotes the model and $s$ the scenario. We assign greater weight to spatial reasoning while retaining distributional fidelity as a grounding constraint. A representative setting is adopted as the default for RSWISE, with detailed validation and sensitivity analyses reported in the experimental section.
@@ 5 Experiment and results
5.1 Implementation Details
We fine-tune BAGEL-7B on action-conditioned remote sensing pairs derived from nearly 4,000 images. Each image is partitioned into overlapping $3{\times}3$ tiles, resulting in 10,080 direction-labeled training instances aligned with the four RSWISE scenarios. Training is conducted on $4{\times}$ H100 (80,GB) GPUs over the course of about 20 hours. For evaluation, we benchmark five models on the RSWISE evaluation set under our protocol, performing inference on $10{\times}$ A100 (80,GB) GPUs with roughly 8,000 runs in total, requiring nearly 80 hours.
@@ 5.1 Implementation Details
Qwen-Image-Edit 46.9 52.1 56.5 57.2 53.2
FLUX.1-Kontext-Dev 40.0 18.7 43.7 41.8 36.1
Step1X-Edit 51.7 17.3 58.5 55.0 45.6
BAGEL 64.3 64.2 62.3 58.7 62.4
@@
Table 2: Performance of different models on RSWISE. Results are reported across four scenarios (General, Flood, Urban, Rural) and their average. Baselines include Qwen-Image-Edit (Wu et al., 2025), FLUX.1-Kontext-Dev (Labs et al., 2025), Step1X-Edit (Liu et al., 2025), and BAGEL (Deng et al., 2025), while RemoteBAGEL is our proposed method.
Figure 3: Comparison of models on FID (lower is better) and GPT (higher is better).
Figure 4: Qualitative comparison of rightward continuations across four benchmark scenarios (general, flood, urban, rural). RemoteBAGEL generates geospatially consistent extrapolations that follow the directional instruction, closely matching the ground truth. By contrast, other models frequently fail to introduce valid extrapolated content or produce semantically inconsistent results (e.g., reversed directions or mismatched structures), leading to large discrepancies from the ground truth.
5.2 Main Results
Our evaluation reveals a clear performance hierarchy among the tested models. RemoteBAGEL substantially outperforms all baselines across four benchmark scenarios, achieving near-optimal scores ($\sim$95) in general and rural settings. This represents a significant advancement over BAGEL (58-64 points), despite BAGEL’s strong multimodal foundations. The performance gap demonstrates that domain-specific adaptation is crucial for remote sensing tasks, as generic vision-language models struggle to capture the spatial coherence and structural patterns inherent in satellite imagery.
5.3 FID vs. GPT Metrics.
Figure 3 highlights the complementary roles of the two metrics. FID, dominated by texture and artifact statistics, offers a reliable measure of visual fidelity and detail preservation, but is less sensitive to semantic plausibility or directional accuracy. GPT-based evaluation, in contrast, directly probes spatial reasoning by scoring continuity, transitions, and compliance with the instructed direction, yet is less attuned to subtle degradations in low-level image quality. This explains why generic editing baselines such as Qwen-Image-Edit and FLUX.1-Kontext-Dev achieve competitive FID values despite failing to introduce meaningful extrapolated content, as reflected in their low GPT scores. At the same time, RemoteBAGEL attains the highest GPT scores across all scenarios (up to 8.98 in rural) while maintaining competitive FID even in challenging settings such as flood. Together, GPT captures the decisive dimension of semantic extrapolation, while FID provides complementary sensitivity to visual quality, sharpening distinctions between strong models like BAGEL and RemoteBAGEL.
5.4 Result Analysis
@@
Qualitative analysis in Figure 4 reveals distinct failure modes across model categories. While BAGEL generates visually plausible outputs, it systematically violates spatial consistency-producing incorrect orientations, missing expected structures (e.g., building clusters), or ignoring geometric regularities in agricultural patterns. Other baselines exhibit more fundamental limitations, generating outputs nearly identical to input tiles with minimal spatial extrapolation, indicating their inability to perform meaningful world modeling. These failures underscore that visual realism alone is insufficient; successful remote sensing continuation requires both semantic understanding and spatial reasoning capabilities that only domain-aware supervision can provide.
@@
Figure 5: FID-based analysis of BAGEL and RemoteBAGEL. (a) Distributional comparison across four scenarios shows that RemoteBAGEL consistently achieves lower FID, with the largest gains in general and rural scenes and the smallest in flood settings. (b) Directional comparison reveals an anisotropic pattern: left and right continuations are easier to model than upward or downward ones, indicating a directional bias in spatial extrapolation.
@@
Figure 5 (b) reports model performance across the four directional prompts and reveals an anisotropic pattern: continuations along one axis (left-right) tend to achieve lower FID, whereas those along the orthogonal axis (up-down) exhibit higher error rates across all scenarios. This asymmetry does not imply a fixed geographic bias—our tiles are not aligned with cardinal directions-but rather reflects differences in local semantic continuity. Horizontal continuations within the image grid often involve more homogeneous structures such as roads or farmland strips, while vertical continuations more frequently cross heterogeneous transitions such as urban-rural boundaries or land-water interfaces. These results suggest that extrapolation errors are amplified when generation requires bridging semantically diverse regions, highlighting a key challenge for spatial reasoning beyond texture fidelity.
Weight analysis.
We validate the RSWISE weighting scheme through a systematic scan of $w_{\text{spatial}}$ (Figure 6). The five diagnostics respectively examine rank correlation, min margin, pareto scatter, score sensitivity, and STD separation. Together they show that rankings remain stable and discriminative power is preserved within the interval $[0.5,0.7]$, supporting the choice of $w_{\text{spatial}}=0.6$ and $w_{\text{fid}}=0.4$. Detailed interpretations of each diagnostic are provided in the Appendix.

### v3

@@ 1 Introduction
World models have emerged as a frontier in artificial intelligence, showing promise across diverse applications such as robotic navigation (wu2023DayDreamer) and autonomous driving (guan2025World). These models aim to learn the compressed latent representations of environments from limited observations and to predict or reason about unobserved states by simulating the underlying dynamics in this latent space (ding2025Understanding). However, most world model studies remain confined to synthetic simulators or constrained scene settings. Synthetic settings lack the complexity and uncertainty of real environments, while constrained scene settings fail to capture reasoning over large spatial structures. As a result, the real-world effectiveness of current world models in spatial reasoning remains largely untested.
@@
Remote sensing provides a uniquely powerful testbed for world models. Satellite and aerial imagery naturally encode “world-level” structures such as urban road networks (yu2023Urban), river systems (tomsett2019Remote), agricultural mosaics (khanal2020Remote), and forest landscapes (fassnacht2024Remote). At the same time, high-impact applications—including flood prediction for disaster response (nguyen2024Solving) and infrastructure forecasting in urban planning (wellmann2020Remote)—require reasoning beyond directly observed regions. Yet, much of remote sensing research has focused on recognition tasks such as classification (li2022DKDFN; temenos2023Interpretable) and semantic segmentation (sun2020BAS\$^4\$Net; zhang2023Semisupervised), leaving the potential of world modeling in this domain unexplored.
@@
This paper bridges these gaps by introducing the first framework for world modeling in remote sensing. We formulate remote sensing world modeling as direction-conditioned spatial extrapolation (defined in the image-grid frame with up, down, left, and right, rather than geographic cardinal directions), where models generate semantically consistent adjacent image tiles given a central observation and directional instruction. As illustrated in Figure 1, this formulation explicitly models the spatial transition as a next-state prediction task, requiring inference over the unobserved world structure.
Figure 1: Illustration of direction-conditioned spatial extrapolation for Remote Sensing World Modeling. Given a central observation $x_{t}$ (the current spatial tile) and a directional instruction $a_{t}$, the world model learns the underlying geospatial structure and predicts the adjacent, previously unobserved tile $x_{t+1}$. Here, the index $t$ denotes spatial progression rather than temporal evolution, aligning our task with the next-state prediction paradigm used in World Models.
@@
However, introducing world models to the remote sensing domain faces a fundamental evaluation challenge rooted in the limitations of existing assessment paradigms. Current evaluation approaches suffer from critical methodological flaws across two distinct failure modes. Distributional fidelity metrics such as Fréchet Inception Distance (FID) (heusel2017GANs) measure statistical realism but ignore whether generated tiles follow spatial instructions. As a result, models could obtain deceptively low FID scores by replicating inputs or producing visually plausible but spatially incoherent imagery. Conversely, large language model–based semantic evaluations such as World Knowledge-Informed Semantic Evaluation (WISE) (niu2025WISE) capture instruction-following semantics but lack quantitative grounding in distributional realism, making it difficult to detect fine-grained issues such as spatial discontinuities or texture mismatches.
@@
To overcome these limitations, we introduce RSWISE (Remote Sensing World-Image Spatial Evaluation), the first evaluation framework designed for remote sensing world models. RSWISE integrates distributional fidelity with spatial reasoning consistency through a dual-dimension approach. Specifically, it employs FID to ensure adherence to real-world satellite statistics and leverages GPT-4o to assess whether generated tiles reveal novel yet geographically plausible regions aligned with directional prompts. As shown in Figure 2, RSWISE reflects better geospatial alignment than FID alone, providing a principled basis for fair comparison and progress tracking.
@@
Building on BAGEL (deng2025emergingpropertiesunifiedmultimodal), a unified multimodal model for generation and understanding, we introduce RemoteBAGEL, the first remote sensing world model specifically designed for direction-conditioned spatial extrapolation. In contrast to prior methods that may yield visually plausible yet spatially inconsistent completions, RemoteBAGEL explicitly couples the generative process with spatial reasoning constraints. It is built around two components: (1) a trajectory-based data construction pipeline that transforms raw satellite imagery into instruction-conditioned continuation tasks, and (2) a reconstruction-driven training framework that enforces geographic continuity and semantic coherence during spatial extrapolation. Extensive experiments demonstrate that RemoteBAGEL consistently outperforms state-of-the-art baselines on RSWISE. In summary, our contributions are threefold:
@@ 2.1 World Models and Benchmarks
World models generally divided into two perspectives: models that aim to understand the world by abstracting its underlying mechanisms, and models that aim to predict the future by simulating possible evolutions of the environment (ding2025Understanding). Early efforts such as World Models (ha2018world) focused on abstracting the external world to gain a deep understanding of its underlying mechanisms, while subsequent work including PlaNet (hafner2019Learning) and the Dreamer family (hafner2020Dream; hafner2022Mastering) introduced recurrent state-space models (RSSMs) that designed to facilitate forward prediction purely within the latent space. More recent advances extend this principle into generative modeling: transformer-based models such as TransDreamer (chen2024TransDreamer) and Genie (bruce2024Genie), diffusion and VAE-driven approaches for scene extrapolation and controllable driving (wang2025DriveDreamer; cai2023DiffDreamer), and the JEPA family (assran2023SelfSupervised) that reframe world models as self-supervised abstraction learners.
Beyond these task-specific designs, emerging unified multimodal models (UMMs) such as BAGEL (deng2025emergingpropertiesunifiedmultimodal) and Qwen-Image-Edit (wu2025qwenimagetechnicalreport) demonstrate the capacity to jointly support both perception (understanding the input image and directional instruction) and generation (synthesizing the new tile). This inherent capability for unified understanding and generation, which includes learning compressed representations of the input and inferring and generating new states, positions UMMs as potential foundation world models. They naturally handle the multimodal input and the complex generation task, overcoming the need for separate, modular components. This work explores how such unified architectures perform when extended to spatial reasoning and world modeling tasks in the remote sensing domain.
Despite this diversity, evaluation remains a central challenge: existing benchmarks such as VBench (ji2024T2VBench), ChronoMagic-Bench (yuan2024ChronoMagicBench), TC-Bench (feng2024TCBench), and WorldModelBench (li2025WorldModelBench) focus on controlled or synthetic scene settings, but they do not explicitly involve spatial continuity or geospatial semantics at the remote sensing scale in real-world contexts. This gap prevents current evaluations from testing how well world models reason over real-world structures such as rivers, roads, or urban–rural transitions, as summarized in Table 1.
@@
VBench ji2024T2VBench 800 Video FID / Human
WorldModelBench li2025WorldModelBench 350 Games / Video Task-specific
ChronoMagic-Bench yuan2024ChronoMagicBench 1649 Video Temporal Consistency
TC-Bench feng2024TCBench 150 Video FID
@@ 2.2 Remote Sensing Models
Remote sensing (RS) acquires Earth observations from satellites and aerial platforms, producing imagery that encodes both spectral variation and large-scale spatial structures across diverse environments (li2019Deep). Much of RS research has traditionally focused on recognition-oriented tasks, including land-cover classification and semantic segmentation (sun2020BAS\$^4\$Net; zhang2023Semisupervised). While recent advances have led to large-scale RS foundation models such as SkySense (guo2024SkySense) and SpectralGPT (hong2024SpectralGPT) that excel at learning transferable representations, these methods seldom attempt spatial continuation or reasoning over large geospatial structures. However, generative AI is an emerging field in RS, with diffusion models being applied to tasks such as super-resolution, cloud removal, and metadata-conditioned image synthesis  (huang2025survey). Across these generative methods-including Text2Earth  (liu2025text2earth) and EMRDM (liu2025effective)-the focus remains strictly on within-tile image synthesis or restoration, rather than modeling the latent structure of the Earth surface or predicting unobserved world states. Therefore, we explore world modeling as a new paradigm for RS, introducing RemoteBAGEL to explicitly couple unified multimodal generation with the spatial reasoning requirements inherent to Earth observation data.
@@ 3 The RSWISE Benchmark
Figure 2: Overview of the RSWISE evaluation framework. RSWISE assesses spatial extrapolation quality along three axes-continuity fidelity, semantic transitions, and directional consistency. These axes are jointly operationalized through two complementary metrics: GPT-4o for spatial reasoning and FID for distributional fidelity. The five examples illustrate how models differ across these axes: some achieve low FID yet fail to produce meaningful directional content, while others show strong spatial reasoning and continuity but incorrect directional consistency. RSWISE integrates both aspects to provide a balanced and geospatially grounded assessment of world modeling performance.
@@ 3.2 Dataset Curation
Figure 3: Start tile is paired with their four cardinal neighbors, yielding evaluation triplets.
To comprehensively evaluate spatial extrapolation under diverse geospatial conditions, we define four representative scenarios that capture complementary challenges in remote sensing world modeling. The first is the general setting of geographic extrapolation, designed to evaluate model capability across diverse scene types rather than within a single environment, thereby enabling a more comprehensive assessment. Beyond this general setting, flood scenarios introduce highly dynamic environmental variations, urban regions emphasize continuity across dense built environments, and rural landscapes highlight consistency within natural and agricultural patterns. These scenarios jointly establish a structured basis for assessing spatial extrapolation across stable, dynamic, structured, and natural contexts.
The benchmark dataset consists of 1,600 curated evaluation instances evenly distributed across these four scenarios: general, flood, urban, and rural. The data are sourced from three publicly available datasets: Sky-SA (zhu2025skysense) for general scenes, FloodNet (rahnemoonfar2021floodnet) for flood events, and LoveDA (wang2021loveda) for urban and rural landscapes.
Although built from three public datasets, the resulting benchmark exhibits substantial diversity: it spans imagery collected across multiple cities and geographically distinct regions, covers a wide range of spatial resolutions from ultra–high-resolution UAV imagery ($\sim$1.5 cm GSD) to satellite imagery at 0.3 m GSD, and includes more than 1,700 distinct semantic categories represented across varied land-cover types. These characteristics ensure that RSWISE evaluates spatial extrapolation under diverse geospatial structures rather than a narrow or dataset-specific distribution.
@@
For each scenario, large satellite images are divided into $3\times 3$ overlapping grids of tiles, where each tile overlaps with its neighbors by approximately $66.7\%$ along both horizontal and vertical directions, ensuring boundary consistency and preserving spatial autocorrelation. Start tiles are paired with their four cardinal neighbors (up, down, left, right), yielding evaluation triplets. Directional instructions are standardized into fixed prompts (e.g., “Look at what is below this picture”) to ensure fairness across models. Filtering criteria include cloud cover thresholds, resolution consistency, and temporal alignment. A quality assurance process-combining automated checks for artifacts, manual inspection of geographic coherence, and balanced sampling—produces 400 instances per category.
@@ 3.3 RSWISE Evaluation Metrics
where $m$ denotes the model and $s$ the scenario. We assign greater weight to spatial reasoning while retaining distributional fidelity as a grounding constraint. A representative setting is adopted as the default for RSWISE. For a detailed validation and sensitivity analysis of the weighting scheme, please refer to Appendix C.
@@ 5 Experiment and results
5.1 Experimental Setting
Dataset Construction.
All training and evaluation data are strictly separated. Although the dataset sources have been described earlier, here we emphasize that the training set and the RSWISE benchmark share no overlapping images or geographic regions. We manually screened all samples to ensure that no training tile originates from the same or even a visually similar location as any test tile, fully preventing data leakage. To construct the training instances, each image is divided into a $3\times 3$ grid, producing nine tiles. Any tile that has at least one adjacent tile is treated as a potential starting tile, and each neighboring tile defines a direction-specific transition. Enumerating all valid start–neighbor combinations yields 24 distinct directional actions per image. Across the 420 selected training images, this procedure results in a total of 10,080 action-conditioned training pairs. And the benchmark utilizes 400 unique satellite images, each divided into a $3\times 3$ grid where the central tile is the starting block. This results in 1,600 test instances, corresponding to the four cardinal direction extrapolation tasks relative to the central starting tile.
Implementation Details.
We fine-tune BAGEL-7B on the 10,080 action-conditioned instances using $4\times$ H100 (80GB) GPUs for approximately 20 hours. For fair comparison, inference for all five models across the full 1,600 benchmark tasks is performed in a strict zero-shot mode, totaling roughly 8,000 runs and requiring about 8 hours of compute using $10\times$ A100 (80GB) GPUs.
Figure 5: Qualitative comparison of rightward continuations across four scenarios (general, flood, urban, rural). RemoteBAGEL produces geospatially consistent extrapolations aligned with the ground truth, whereas other models often generate invalid or semantically inconsistent content.
5.2 Main Results
Our evaluation reveals a clear performance hierarchy among the tested models. RemoteBAGEL substantially outperforms all baselines across four benchmark scenarios, achieving near-optimal scores ($\sim$95) in general and rural settings. This represents a significant advancement over BAGEL (58-64 points), despite BAGEL’s strong multimodal foundations. The performance gap demonstrates that domain-specific adaptation is crucial for remote sensing tasks, as generic vision-language models struggle to capture the spatial coherence and structural patterns inherent in satellite imagery. Furthermore, an Out-of-Distribution (OOD) test demonstrates RemoteBAGEL’s robust generalization capability in unseen hurricane scenarios, with detailed results provided in Appendix G.
5.3 Result Analysis
@@
Qwen-Image-Edit (wu2025qwenimagetechnicalreport) 46.9 52.1 56.5 57.2 53.2
FLUX.1-Kontext-Dev (labs2025flux1kontextflowmatching) 40.0 18.7 43.7 41.8 36.1
Step1X-Edit (liu2025step1xeditpracticalframeworkgeneral) 51.7 17.3 58.5 55.0 45.6
BAGEL (deng2025emergingpropertiesunifiedmultimodal) 64.3 64.2 62.3 58.7 62.4
@@
Table 2: Performance of different models on RSWISE. Results are reported across four scenarios (General, Flood, Urban, Rural) and their average.
Figure 6: FID-based analysis of BAGEL and RemoteBAGEL. (a) Distributional results across four scenarios show consistently lower FID for RemoteBAGEL, with the largest gains in general and rural settings. (b) Directional results reveal an anisotropic pattern, where left/right continuations are easier than up/down, indicating directional bias in spatial extrapolation.
@@
Qualitative analysis in Figure 5 reveals distinct failure modes across model categories. While BAGEL generates visually plausible outputs, it occasionally violates spatial consistency-producing incorrect orientations or ignoring geometric regularities. However, compared to baselines, its potential for reasoning is evident. Crucially, although baseline approaches also utilize unified architectures, they fail to effectively propagate the commonsense knowledge from their frozen large language model (LLM) backbones into the visual generation process. Consequently, they operate primarily as appearance editors, generating outputs that mimic input textures but lack genuine spatial extrapolation. BAGEL partially overcomes this by leveraging better multimodal alignment, yet it still falters due to a lack of domain-specific grounding; it struggles to map abstract direction tokens to the complex spatial semantics of satellite imagery. In contrast, the superior performance of RemoteBAGEL demonstrates that our fine-tuning strategy successfully bridges this gap. It allows the model to internalize geospatial priors and effectively generalize the backbone’s inherent reasoning capabilities to the task of direction-aware spatial world modeling.
@@
Figure 6 (b) illustrates an anisotropic performance pattern where horizontal continuations consistently outperform vertical ones. We attribute this asymmetry to the interplay between physical imaging factors and the model’s inductive biases. Physically, variations in solar azimuth and the polar orbital paths of satellites introduce greater radiometric and geometric inconsistencies along the vertical axis, creating naturally harder extrapolation targets compared to the more stable illumination in horizontal directions. From a modeling perspective, this anisotropy likely reflects the distribution of spatial priors in the pre-trained LLM backbone. In general multimodal corpora, horizontal relationships (e.g., reading order, panoramic layouts) are often represented with stronger continuity and concrete logical links than vertical ones. Consequently, the direction tokens for left/right may activate more robust spatial reasoning patterns within the model than up/down tokens. This combination of inherent domain challenges (lighting/orbit) and model-level inductive bias (weaker vertical priors) results in the observed performance gap.
5.4 Ablation Studies
[…]
Inference Phase
Prompting Strategy
No prompt 52.3 215.4 4.10
Cardinal (N/S/E/W) 72.1 201.5 7.20
Grid-aligned (Ours) 88.8 196.0 8.86
[2pt/2pt] Overlap Ratio
0% 78.5 235.8 8.10
33% 82.1 215.5 8.45
66.7% (Ours) 88.8 196.0 8.86
Training Phase
Directional Conditioning
No direction token 58.7 198.5 5.10
With token (Ours) 88.8 196.0 8.86
[2pt/2pt] Overlap Strategy
0% 74.6 235.1 6.80
33% 80.2 210.4 7.95
66.7% (Ours) 88.8 196.0 8.86
Table 3: Ablation studies on Inference and Training configurations.
We validate key design choices via controlled ablations (details in Appendix F). Results in Table  3 highlight three findings:
Prompt Formulation. Grid-aligned instructions (“up/down”) significantly outperform cardinal directions (“north/south”). The model exhibits poor responsiveness to geographic terms, failing to accurately execute the specific directional commands compared to image-relative coordinates.
Spatial Overlap. High spatial overlap (66.7%) is fundamental for maintaining geospatial coherence. Reducing overlap disrupts boundary continuity during inference and prevents the model from learning valid cross-boundary transitions during training.
Directional Conditioning. Explicit direction tokens are essential. Removing them leads to a drastic performance drop to 58.7, resulting in random generation where the model fails to explicitly generate content according to the instructions.
5.5 Validity of GPT-Based Evaluation
We confirmed the scientific rigor of our GPT-4o metric via two studies: multi-run stability (mean standard deviation 0.026) and human agreement (Spearman $\rho=0.72$). These results establish the GPT-based metric as a reliable and expert-aligned semantic evaluator, essential for distinguishing genuine spatial reasoning from visual plagiarism. Detailed statistics are provided in Appendix E.
Figure 7: Continuous multi-step extrapolation results. Compared with the ground truth (GT), the generated results remain plausible in the first step but gradually deviate as extrapolation continues.
5.6 Multi-Step Extrapolation
Figure 7 illustrates the performance of the model under continuous multi-step extrapolation. RemoteBAGEL maintains high visual and structural consistency with the ground truth (GT) during the early steps (Step 1). However, as extrapolation proceeds (Step 3–4), noticeable drifts emerge—boundaries blur, textures repeat, and geometric structures gradually distort. This degradation primarily results from cumulative error propagation: at each step, the model conditions on its own generated output, causing small local deviations to amplify over time and drift from the true spatial layout. The absence of global constraints across successive steps further weakens long-range coherence, leading to compounded structural inconsistencies. Overall, RemoteBAGEL exhibits strong short-range continuation ability but remains limited by recursive error accumulation and context drift in extended generation.